OpenStudy (anonymous):

What is the domain and range of the function y=7/2x-10 ?

OpenStudy (anonymous):

you mean y=7/(2x-10) or y=(7/2x)- 10

OpenStudy (anonymous):

is it \[y=\frac{7}{2x-10}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you cannot divide by 0 so x cannot be 5

OpenStudy (anonymous):

so domain is all real numbers except 5 range: ratio can never be 0 because for a fraction to be 0 the numerator has to be 0 and this numerator is 7

OpenStudy (anonymous):

so range is all numbers except 0

OpenStudy (anonymous):

Oh ok so how would i write that as interval notation?

OpenStudy (anonymous):

doamin all real numbers except 5 ,

OpenStudy (anonymous):

it is stupid to write all numbers except 5 as an interval because it is easier to say it. but if you have to, you can write \[(-\infty,5)\cup(5,\infty)\] which a rather silly way to say "exclude 5"

OpenStudy (anonymous):

LOl alrity thank you.